# Compositional Generative Model of Unbounded 4D Cities

### v2

@@ 5.5 Applications
Localized Editing. Benefiting from the compositional architecture, CityDreamer4D allows for localized editing on building and vehicle instances. In Fig. 10(a) and 10(c), vehicle positions and styles can be independently modified without affecting other scene elements. Similarly, as shown in Fig. 10(b) and 10(d), building appearances adapt seamlessly to varying heights while maintaining a consistent style. This capability facilitates customized scene refinement in post-production.
@@ 5.6 Discussions
Relighting. In CityDreamer4D, the generation of background stuff and instances is deliberately decoupled, offering two key benefits: (1) Simplified learning for building instances, vehicle instances, and background stuff, and (2) Enabling localized editing of building and vehicle instances. This approach can be viewed as an inverse rendering process, where CityDreamer4D generates the albedo, normals, and depth of urban scenes. Lighting and shading effects are then computed based on the given lighting conditions. As shown in Fig. 13, the shading effects are divided into two components: Lambertian shading and shadow mapping. Lambertian shading accounts for the light direction and surface normal, resulting in uniform lighting across all directions, as shown in Fig. 13(a). Shadow mapping considers light visibility, enabling the simulation of shadows and occlusion from other objects in the scene, as shown in Fig. 13(b). The final relighting effects, with the camera positioned on the left side of the scene, are presented in Fig. 13(c).

### v4

@@ 5.5 Applications
Localized Editing. Benefiting from the compositional architecture, CityDreamer4D allows for localized editing on building and vehicle instances. In Fig. 10a and 10c, vehicle positions and styles can be independently modified without affecting other scene elements. Similarly, as shown in Fig. 10b and 10d, building appearances adapt seamlessly to varying heights while maintaining a consistent style. This capability facilitates customized scene refinement in post-production.
@@
TABLE VI: Visual Language Navigation (VLN) results in generated 4D cities. Metrics include: PL (Path Length, scaled by 1/10, in meters), SR (Success Rate, %), SPL (Success rate weighted by normalized inverse Path Length), and RT (Reset Times). Note that each trajectory is executed five times and is considered a failure if all attempts result in resets.
[…]
Human (5 participants) - 20.73 100.0 85.87 0.00
Gemini 2.5 Pro [117] - 9.32 12.40 4.43 0.45
GPT-4o [118] - 8.97 36.00 17.32 0.11
SAIL-VL 1.6 [119] 8.33 14.56 23.40 7.63 0.28
Ovis2 [120] 8.94 13.96 17.00 5.01 0.35
Qwen2.5-VL [121] 8.29 5.01 15.00 7.01 0.37
Ola [122] 8.88 9.15 18.00 8.30 0.32
InternVL3 [123] 7.94 9.02 25.60 12.66 0.23
Visual Language Navigation. To evaluate the practicality of our generated 4D urban environments, we conduct experiments on Visual Language Navigation, where an embodied agent navigates based on natural language instructions. Specifically, the agent operates within scenes generated by CityDreamer4D, pretrained on the CityTopia dataset. We manually annotate a test set of 100 instruction-trajectory pairs, each guiding the agent to a distinct landmark in the generated scenes. Following the protocol in GRUtopia [124], the agent receives its current image observation and a language prompt as input to a vision-language model (VLM), which selects one of 12 discrete actions: move forward or diagonally (2/4/6 meters), turn left/right (45°), or stop. This process continues iteratively until the model outputs “stop”. We additionally collect a human performance reference from 5 volunteers, who complete the full test set following the same navigation protocol as the VLMs. We adopt the zero-shot evaluation setting from GRUtopia, using pre-trained VLMs without task-specific fine-tuning. We evaluate the latest state-of-the-art VLMs, including open-source models such as SAIL-VL 1.6 [119], Ovis2 [120], Qwen2.5-VL [121], Ola [122], and InternVL3 [123], as well as closed-source models like Gemini 2.5 Pro [117] and GPT-4o [118]. Navigation performance is evaluated using four standard metrics: success rate (SR), path length (PL), success rate weighted by normalized inverse path length (SPL), and the number of resets due to occlusions (RT). As shown in Table VI, VLMs struggle with spatial reasoning in 4D cities, as indicated by low SR and SPL scores. GPT-4o performs best, followed by InternVL3. For comparison, the human reference achieves an SR of 100.0% and an SPL of 85.87%, indicating a substantial gap between human and model performance. The average human path length is slightly higher than that of VLMs, mainly because the latter often stop before actually reaching the destination. These results, along with recent findings [124, 125], underscore the difficulty of grounding spatial instructions in complex urban environments and suggest that generated 4D cities can serve as valuable benchmarks for evaluating the spatial reasoning and navigation capabilities of VLMs.
@@
Figure 14: Pedestrians in the generated 4D cities. The three consecutive frames illustrate a group of pedestrians crossing the street.
@@ 5.6 Discussions
Relighting. In CityDreamer4D, the generation of background stuff and instances is deliberately decoupled, offering two key benefits: (1) Simplified learning for building instances, vehicle instances, and background stuff, and (2) Enabling localized editing of building and vehicle instances. This approach can be viewed as an inverse rendering process, where CityDreamer4D generates the albedo, normals, and depth of urban scenes. Lighting and shading effects are then computed based on the given lighting conditions. As shown in Fig. 13, the shading effects are divided into two components: Lambertian shading and shadow mapping. Lambertian shading accounts for the light direction and surface normal, resulting in uniform lighting across all directions, as shown in Fig. 13a. Shadow mapping considers light visibility, enabling the simulation of shadows and occlusion from other objects in the scene, as shown in Fig. 13b. The final relighting effects, with the camera positioned on the left side of the scene, are presented in Fig. 13c.
Diverse Agents Support. To explore the potential for supporting more diverse agents in CityDreamer4D, we conduct a preliminary experiment integrating pedestrians into the generated scenes. We use MoMask [126] to synthesize motion, retarget it to 3D human avatars, and render the animated pedestrians into our generated 4D environments. As shown in Figure 14, the resulting animation demonstrates coherent pedestrian behavior such as street crossing. This highlights the feasibility of extending our framework to support richer multi-agent simulations beyond vehicles.